OpenStudy (anonymous):

which polynomial represents (x+5)(3x-2)

OpenStudy (amistre64):

they tend to teach a method called FOIL .... this is where you would apply it

OpenStudy (amistre64):

you could also do a staked method x +5 3x -2 multiply down the columns, and a crisscross pattern add all the results

hero (hero):

(x+5)(3x-2) = x(3x - 2) + 5(3x - 2) = 3x^2 - 2x + 15x - 10 = ?